An investment of $3,000 is made at an annual simple interest of 5%. How much additional money must be invested at an annual simple interest rate of 8%, so that the total annual interest earned is 7.5 of the original amount that you invested?

Answer by Theo(13342)

the original amount you invested is 3000.

7.5 * 3000 = 22,500

5% of 3000 = .05 * 3000 = 150.

the additional amount invested at 8% must be able to give you 22,000 - 150 = 22,350.

let x equal that amount.

formula becomes x * .08 = 22,350.

divide both sides of this equation by .08 to get x = 22,350/.08 = 279,375.

that's how much you would need to invest at 8% in order to get interest of 22,350.

the interest of 22,350 plus the interest of 150 is equal to 22,500 which is 7.5 * the original investment of 3000.

i'm not exactly sure this is what you meant, but, based on what i think you meant, this is the solution.